OpenStudy (anonymous):

What is the slope for (8,-2) and (7,-2)

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

i got -4/1

OpenStudy (anonymous):

I feel like that is not right

hartnn (hartnn):

what is y1 and y2 ??

OpenStudy (anonymous):

-2,-2

OpenStudy (anonymous):

Take a look at this link it WILL help you! http://cs.selu.edu/~rbyrd/math/slope/

hartnn (hartnn):

\(y_1-y_2 = -2-(-2) = -2+2 =0\)

hartnn (hartnn):

your numerator itself is 0! so what would be the slope ?

OpenStudy (anonymous):

so 0/1 automatically is a slope of 0 because of the numerator!

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (8,-2)...x1 = 8 and y1 = -2 (7,-2)...x2 = 7 and y2 = -2 now we sub slope = (-2 - (-2) / (7 - 8) slope = (-2 + 2) / -1 slope = 0/-1 if the 0 is in the numerator of the fraction, then it is a horizontal line with a slope of 0.

OpenStudy (anonymous):

gotcha. Thank you so much. this was very helpful

OpenStudy (texaschic101):

if 0 would have been in the denominator, then it would have been a vertical line with a slope undefined

hartnn (hartnn):

thats correct :) slope = 0

hartnn (hartnn):

bdw, since you're new here \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)